OpenStudy (anonymous):

prove that the identity: sin(θ+π/3)+cos(θ+π/6)=√3 cosθ

OpenStudy (anonymous):

You accidentally an extra word.

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

"Prove the identity" not "Prove that the identity"

OpenStudy (tyteen4a03):

@SmoothMath Not very productive ;) Remember our angle addition and subtraction formula: \(cos(A + B) = cos(A)cos(B) - sin(A)sin(B)\)

OpenStudy (shubhamsrg):

cos(θ+π/6) = cos(θ+π/2 - π/3) = sin( π/3 - θ) you may now use sinA + sin B identity to directly reach an answer.